OpenStudy (anonymous):

If u = x^2 cos 2x find d/dx cosh u If u = x^2 cos 2x find d/dx cosh u @Mathematics

OpenStudy (anonymous):

(d/dx)cosh(u)=(du/dx)*sinh(u) =(2x*cos(2x)-2x^2*sin(2x))*sinh(x^2*cos(2x))

OpenStudy (anonymous):

The question is to find \(\frac{d}{dx} \cosh u\), where u is some function of \(x\). You need to apply the chain rule here: \(\frac{d}{dx}\cosh u=\sinh u.u'=(-2x^2\sin(2x)+2x\cos(2x))\sinh( x^2 \cos 2x).\)

OpenStudy (anonymous):

You can simplify it a little bit if you want: \(\large{\frac{d}{dx}\cosh u=2x(\cos(2x)-x\sin(2x))\sinh(x^2\cos(2x))}\).